A Practice Based M Level Course

The idea that a degree was formally taken by the applicant showing himself competent for it, may be well illustrated from the quaint ceremony of admitting a Master in Grammar at Cambridge, as described by the Elizabethan Esquire Bedel, Mr. Stokys: ’The Bedel in Arts shall bring the Master in Grammar to the Vice-Chancellor, delivering him a palmer with a rod, which the Vice-Chancellor shall give to the said Master in Grammar, and so create him Master. Then shall the Bedel purvey for every Master in Grammar a shrewd boy, whom the Master in Grammar shall beat openly in the Schools, and he shall give the boy a groat for his labour, and another groat to him that provideth the rod and the palmer. And thus endeth the Act in that faculty.’

[See ‘The Oxford Degree Ceremony’, J. Wells OUP 1906 on Guttenberg]‘